# Thermodynamic speed limit for non-adiabatic work and its classical-quantum decomposition

### v1

@@ IV.2 Thermodynamic speed limits for non-adiabatic classical work
$\dot{W}_{\rm cl}-\dot{W}_{\rm ad}=\mathrm{Tr}[\dot{H}_{t}^{\rm cl}\mathcal{D}_{t}^{+}[\mathcal{D}_{t}[\rho_{t}^{\rm diag}]]],$ (24)
where $\mathcal{D}_{t}^{+}$ is the Drazin inverse of $\mathcal{D}_{t}$. By following a derivation similar to that for Eq. (14), which is detailed in Appendix C, the non-adiabatic classical work $\dot{W}_{\rm cl}-\dot{W}_{\rm ad}$ is bounded as
$|\dot{W}_{\rm cl}-\dot{W}_{\rm ad}|\leq\sqrt{2A(\rho_{t}^{\rm diag})\dot{\sigma}_{\rm cl}}=:\mathcal{B}_{2}^{\rm cl},$ (25)
@@
$\dot{\sigma}_{\rm cl}:=\dot{S}_{\rm diag}-\beta\dot{Q}\geq 0$ (26)
quantifies the incoherent part of the entropy production, with $S_{\rm diag}:=S(\rho_{t}^{\rm diag})$ being the diagonal entropy. We can further show that [59, 60, 40] (see Eq. (82))
@@
$A(\rho_{t}^{\rm diag})=\frac{1}{2}\sum_{n}\langle\epsilon_{n}|\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]|\epsilon_{n}\rangle^{2}\langle\epsilon_{n}|\mathcal{D}_{t}^{\rm sym}[\rho_{t}^{\rm diag}]|\epsilon_{n}\rangle.$ (28)
This expression follows from the definition of $A(\rho_{t})$ in Eq. (16), by noting that the diagonal basis of $\rho_{t}^{\rm diag}$ is given by $|\epsilon_{n}\rangle$, and that $\langle\epsilon_{n}|\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]|\epsilon_{n}\rangle=\langle\epsilon_{n}|\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]|\epsilon_{n}\rangle$, where $\tilde{\mathcal{D}}_{t}^{+}$ is the adjoint of $\mathcal{D}_{t}^{+}$.
@@ IV.3 Thermodynamic speed limits for non-adiabatic quantum work
$\dot{W}_{\rm qm}=\mathrm{Tr}[\dot{H}_{t}^{\rm qm}\mathcal{H}_{t}^{+}[\mathcal{H}_{t}[\rho_{t}]]],$ (29)
where $\mathcal{H}_{t}^{+}$ is the Drazin inverse of $\mathcal{H}_{t}$ defined in Eq. (43). Now, $\dot{W}_{\rm qm}$ can be bounded as (see Appendix C for the proof):
@@
is the quantum Fisher information that quantifies the coherence of the density operator $\delta\rho_{t}^{\rm coh}$, and
@@
is the counter-diabatic Hamiltonian that is used as a control Hamiltonian to guide the system along the instantaneous energy eigenstate [61].
@@
Note that the connections between the variance of $H_{\rm cd}$ and quantum speed limits have been discussed in several studies, including Refs. [62, 63, 24]. We further note that from Eq. (47), we have
@@ IV.4 Classical-quantum decomposition of the thermodynamic speed limit
$|\dot{W}-\dot{W}_{\rm ad}|\leq\underbrace{\frac{1}{\hbar}\sqrt{\mathcal{F}_{\rho_{t}}(H_{t})V_{\rho_{t}}(H_{\rm cd})}}_{\text{quantum}}+\underbrace{\sqrt{2A(\rho_{t}^{\rm diag})\dot{\sigma}_{\rm cl}}}_{\text{classical}}=\mathcal{B}_{2}^{\rm qm}+\mathcal{B}_{2}^{\rm cl}.$ (34)
These classical-quantum decomposition of the thermodynamic speed limits for non-adiabatic work [Eqs. (25), (30), and (34)] are the second main result of this paper.
@@
$|\dot{W}-\dot{W}_{\rm ad}|\leq\min\left\{\mathcal{B}_{1},\ \mathcal{B}_{2}^{\rm qm}+\mathcal{B}_{2}^{\rm cl}\right\},$ (35)
@@ V.1 Classical drive
It should also be noted that when the initial density operator satisfies $\rho(0)=\rho^{\rm diag}(0)$, we have $\rho(t)=\rho^{\rm diag}(t)$ at later times $t$. By assuming this initial condition, we have $g_{tt}^{\rm QF}=0$ and $\dot{\sigma}_{\rm cl}=\dot{\sigma}$. Note that this setup is essentially equivalent to considering classical stochastic systems, and the first main result (14) also reduces to Eq. (36).
@@ VI Example: two-level system
Figure 1: Numerical calculation of the thermodynamic speed limits for non-adiabatic work for two-level systems. Red, green, and blue solid curves plot the non-adiabatic works $|\dot{W}-\dot{W}_{\rm ad}|$, $|\dot{W}_{\rm qm}|$, and $|\dot{W}_{\rm cl}-\dot{W}_{\rm ad}|$, respectively. Black and red dashed curves plot the non-adiabatic work bounds $\mathcal{B}_{1}$ (14) and $\mathcal{B}_{2}^{\rm qm}+\mathcal{B}_{2}^{\rm cl}$ (34), respectively. Blue and green dashed curves are the classical and quantum bounds $\mathcal{B}_{2}^{\rm cl}$ (25) and $\mathcal{B}_{2}^{\rm qm}$ (30), respectively. (a) General drive. Both classical and quantum terms contribute to the non-adiabatic work and the thermodynamic speed limit (34). (b) Classical drive $\dot{H}_{t}=\dot{H}_{t}^{\rm cl}$. Note that quantum contributions vanish and the red and blue curves become identical (see also Sec. V.1). (c) Quantum drive $\dot{H}_{t}=\dot{H}_{t}^{\rm qm}$. Note that classical contributions vanish and the red and green curves become identical (see also Sec. V.2). The parameters are $\gamma_{0}=\beta=1$, $\langle\epsilon_{+}|\rho(0)|\epsilon_{+}\rangle=0.3,\ \langle\epsilon_{+}|\rho(0)|\epsilon_{-}\rangle=0.2+0.1i$.
@@
$H(t)=\frac{\Delta_{t}}{2}\sigma_{x}+\frac{q_{t}}{2}\sigma_{z},$ (38)
where $\Delta_{t}$ and $q_{t}$ denote the time-dependence of the external driving field ($\lambda_{t}=(\Delta_{t},q_{t})$) and $\sigma_{x}$, $\sigma_{z}$ are the Pauli-X-operator and Pauli-Z-operator, respectively. The energy eigenvalues are given by $\epsilon_{\pm}=\pm\omega_{t}/2$, where we define the energy difference as $\omega_{t}:=\sqrt{\Delta_{t}^{2}+\lambda_{t}^{2}}$. The energy eigenstates are given by $|\epsilon_{+}\rangle=\cos\theta_{t}|e\rangle+\sin\theta_{t}|g\rangle$ and $|\epsilon_{-}\rangle=\sin\theta_{t}|e\rangle-\cos\theta_{t}|g\rangle$, where $\theta_{t}:=(1/2)\text{cot}^{-1}(q_{t}/\Delta_{t})$, and $|e\rangle$ and $|g\rangle$ are the eigenstates of $\sigma_{z}$, i.e., $\sigma_{z}|e\rangle=|e\rangle$ and $\sigma_{z}|g\rangle=-|g\rangle$. The time-dependent Hamiltonian (38) can be realized, for example, in the charge qubit configuration, where $q_{t}$ can be varied by changing the applied magnetic flux, while $\Delta_{t}$ can be varied by changing the applied gate voltage [64].
@@
$\partial_{t}\rho_{t}=-i[H_{t},\rho_{t}]+\gamma_{\downarrow}\left(\sigma_{t}^{-}\rho\sigma_{t}^{+}-\frac{1}{2}\left\{\sigma_{t}^{+}\sigma_{t}^{-},\rho_{t}\right\}\right)+\gamma_{\uparrow}\left(\sigma_{t}^{+}\rho\sigma_{t}^{-}-\frac{1}{2}\left\{\sigma_{t}^{-}\sigma_{t}^{+},\rho_{t}\right\}\right),$ (39)
where $\sigma_{t}^{+}=\left.\left|\epsilon_{+}\right\rangle\right.\left.\left\langle\epsilon_{-}\right.\right|$ and $\sigma_{t}^{-}=\left.\left|\epsilon_{-}\right\rangle\right.\left.\left\langle\epsilon_{+}\right.\right|$ are the raising and lowering operators between energy eigenstates, and $\gamma_{\downarrow}:=\gamma_{0}(N(\omega_{t})+1)$ and $\gamma_{\uparrow}:=\gamma_{0}N(\omega_{t})$ with $N(\omega_{t}):=1/(e^{\beta\omega_{t}}-1)$ are the transition rates satisfying the detailed balance condition $\gamma_{\downarrow}/\gamma_{\uparrow}=e^{\beta\omega_{t}}$.
@@ VII Conclusion
An interesting future direction is to further investigate the classical-quantum decomposition of non-adiabatic work and analyze the role of quantum coherence in thermodynamics. In the quasi-adiabatic regime, it has been reported in Ref. [18] that generation of quantum coherence violates the fluctuation-dissipation relation. It has also been observed in various models that generation of quantum coherence is typically detrimental to the performance of heat engines [64]. However, a general understanding on the effect of having coherence in the non-adiabatic work extraction processes is not complete and deserves further studies. Understanding the role of coherence in thermodynamics is anticipated to lead to obtaining the design principle of fast and efficient energy conversion and information processing devices.
It is also interesting to investigate further the geometric and thermodynamic properties of the obtained bound for nonequilibrium states. Extending the concepts of geometric quantities defined for adiabatic or equilibrium states and their optimization methods [15, 16, 17, 18, 19] to nonequilibrium setting is another important future direction.
@@ Appendix A Properties of the Drazin pseudo inverse and its adjoint
Next, let us consider the Drazin inverse of $\mathcal{H}_{t}$ (2). We first note that the steady state of $\mathcal{H}_{t}$ is no longer unique, since any $\rho_{t}^{\rm diag}$ satisfies $\mathcal{H}_{t}[\rho_{t}^{\rm diag}]=0$. For given $\rho_{t}$, the steady-state of $\mathcal{H}_{t}$ reads
$\lim_{\nu\rightarrow\infty}e^{\nu\mathcal{H}_{t}}\rho_{t}=\sum_{n}\langle\epsilon_{n}|\rho_{t}|\epsilon_{n}\rangle|\epsilon_{n}\rangle\langle\epsilon_{n}|=\rho_{t}^{\rm diag},$ (42)
@@
$\mathcal{H}_{t}^{+}[\rho_{t}^{\rm diag}]=\mathrm{Tr}\left[\mathcal{H}_{t}^{+}[X]\right]=0.$ (45)

### v2

@@ IV.2 Thermodynamic speed limits for non-adiabatic classical work
$\dot{W}_{\rm cl}-\dot{W}_{\rm ad}=\mathrm{Tr}[\dot{H}_{t}^{\rm cl}(\rho_{t}^{\rm cl}-\rho_{t}^{\rm eq})]=\mathrm{Tr}[\dot{H}_{t}^{\rm cl}\mathcal{D}_{t}^{+}[\mathcal{D}_{t}[\rho_{t}^{\rm cl}]]],$ (24)
where $\mathcal{D}_{t}^{+}$ is the Drazin inverse of $\mathcal{D}_{t}$, and we have used the relation $\rho_{t}^{\rm cl}-\rho_{t}^{\rm eq}=\mathcal{D}_{t}^{+}[\mathcal{D}_{t}[\rho_{t}^{\rm cl}]]$, which follows from Eq. (45). By following a derivation similar to that for Eq. (14), which is detailed in Appendix C, the non-adiabatic classical work $\dot{W}_{\rm cl}-\dot{W}_{\rm ad}$ is bounded as
$|\dot{W}_{\rm cl}-\dot{W}_{\rm ad}|\leq\sqrt{2A(\rho_{t}^{\rm cl})\dot{\sigma}_{\rm cl}}=:\mathcal{B}_{2}^{\rm cl},$ (25)
@@
$\dot{\sigma}_{\rm cl}:=\dot{S}_{\rm cl}-\beta\dot{Q}\geq 0$ (26)
quantifies the incoherent part of the entropy production, with $S_{\rm cl}:=S(\rho_{t}^{\rm cl})$ being the diagonal entropy. We can further show that Santos et al. (2019); Landi and Paternostro (2021); Tajima and Funo (2021) (see Eq. (86))
@@
$A(\rho_{t}^{\rm cl})=\frac{1}{2}\sum_{n}\langle\epsilon_{n}|\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]|\epsilon_{n}\rangle^{2}\langle\epsilon_{n}|\mathcal{D}_{t}^{\rm sym}[\rho_{t}^{\rm cl}]|\epsilon_{n}\rangle.$ (28)
This expression follows from the definition of $A(\rho_{t})$ in Eq. (16), by noting that the diagonal basis of $\rho_{t}^{\rm cl}$ is given by $|\epsilon_{n}\rangle$, and that $\langle\epsilon_{n}|\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]|\epsilon_{n}\rangle=\langle\epsilon_{n}|\tilde{\mathcal{D}}_{t}^{+}[\dot{H}_{t}^{\rm cl}]|\epsilon_{n}\rangle$, where $\tilde{\mathcal{D}}_{t}^{+}$ is the adjoint of $\mathcal{D}_{t}^{+}$.
@@ IV.3 Thermodynamic speed limits for non-adiabatic quantum work
$\dot{W}_{\rm qm}=\mathrm{Tr}[\dot{H}_{t}^{\rm qm}\rho_{t}^{\rm qm}]=\mathrm{Tr}[\dot{H}_{t}^{\rm qm}\mathcal{H}_{t}^{+}[\mathcal{H}_{t}[\rho_{t}]]],$ (29)
where $\mathcal{H}_{t}^{+}$ is the Drazin inverse of $\mathcal{H}_{t}$ defined in Eq. (47), and we have used the relation $\rho_{t}^{\rm qm}=\mathcal{H}_{t}^{+}[\mathcal{H}_{t}[\rho_{t}]]$, which follows from Eq. (48). Now, $\dot{W}_{\rm qm}$ can be bounded as (see Appendix C for the proof):
@@
is the quantum Fisher information that quantifies the coherence of the density operator $\rho_{t}^{\rm qm}$, and
@@
is the counter-diabatic Hamiltonian that is used as a control Hamiltonian to guide the system along the instantaneous energy eigenstate Guéry-Odelin et al. (2019). Note that $V_{\rho_{t}}(H_{\rm cd})$ that appears in the bound $\mathcal{B}_{2}^{\rm qm}$ is the energy fluctuation of the control Hamiltonian $H_{\rm cd}$, which quantifies the speed of controlling non-adiabatic time-evolution in isolated Funo et al. (2017); Bukov et al. (2019) and quasi-adiabatic open systems Funo et al. (2019). Moreover, $V_{\rho_{t}}(H_{\rm cd})$ sets an upper bound on the work fluctuation in isolated systems controlled by $H_{\rm cd}$ Funo et al. (2017).
@@
We further note that from Eq. (51), we have
@@ IV.4 Classical-quantum decomposition of the thermodynamic speed limit
$|\dot{W}-\dot{W}_{\rm ad}|\leq\underbrace{\frac{1}{\hbar}\sqrt{\mathcal{F}_{\rho_{t}}(H_{t})V_{\rho_{t}}(H_{\rm cd})}}_{\text{quantum}}+\underbrace{\sqrt{2A(\rho_{t}^{\rm cl})\dot{\sigma}_{\rm cl}}}_{\text{classical}}=\mathcal{B}_{2}^{\rm qm}+\mathcal{B}_{2}^{\rm cl}.$ (34)
These classical-quantum decomposition of the thermodynamic speed limits for non-adiabatic work [Eqs. (25), (30), and (34)] are the second main result of this paper. We again note that the splitting of the bound into $\mathcal{B}_{2}^{\rm qm}$ and $\mathcal{B}_{2}^{\rm cl}$ is based on the decomposition with respect to the non-diagonal and diagonal contributions using the basis of the Hamiltonian $H_{t}$, respectively.
@@
$|\dot{W}-\dot{W}_{\rm ad}|\leq\min\left\{\mathcal{B}_{1}^{\rm nd}+\mathcal{B}_{1}^{\rm d},\ \mathcal{B}_{2}^{\rm qm}+\mathcal{B}_{2}^{\rm cl}\right\},$ (35)
@@ V.1 Classical drive
It should also be noted that when the initial density operator satisfies $\rho_{0}=\rho^{\rm cl}_{0}$, we have $\rho_{t}=\rho^{\rm cl}_{t}$ at later times $t$. By assuming this initial condition, we have $g_{tt}^{\rm QF}=0$ and $\dot{\sigma}_{\rm cl}=\dot{\sigma}$. Note that this setup is essentially equivalent to considering classical stochastic systems, and the first main result (14) also reduces to Eq. (36).
@@ VI Example: two-level system
Figure 1: Numerical calculation of the thermodynamic speed limits for non-adiabatic work for two-level systems. Red, green, and blue solid curves plot the non-adiabatic works $|\dot{W}-\dot{W}_{\rm ad}|$, $|\dot{W}_{\rm qm}|$, and $|\dot{W}_{\rm cl}-\dot{W}_{\rm ad}|$, respectively. Black and red dashed curves plot the non-adiabatic work bounds $\mathcal{B}_{1}^{\rm nd}+\mathcal{B}_{1}^{\rm d}$ (14) and $\mathcal{B}_{2}^{\rm qm}+\mathcal{B}_{2}^{\rm cl}$ (34), respectively. Blue and green dashed curves are the classical and quantum bounds $\mathcal{B}_{2}^{\rm cl}$ (25) and $\mathcal{B}_{2}^{\rm qm}$ (30), respectively. (a) General drive. Both classical and quantum terms contribute to the non-adiabatic work and the thermodynamic speed limit (34). (b) Classical drive $\dot{H}_{t}=\dot{H}_{t}^{\rm cl}$. Note that quantum contributions vanish and the red and blue curves become identical (see also Sec. V.1). (c) Quantum drive $\dot{H}_{t}=\dot{H}_{t}^{\rm qm}$. Note that classical contributions vanish and the red and green curves become identical (see also Sec. V.2). The parameters are $\gamma_{0}=\beta=1$, $p_{+}(0)=0.3,\ \rho_{+-}(0)=0.2+0.1i$.
@@
$H_{t}=\frac{\Delta_{t}}{2}\sigma_{x}+\frac{q_{t}}{2}\sigma_{z},$ (38)
where $\Delta_{t}$ and $q_{t}$ denote the time-dependence of the external driving field ($\lambda_{t}=(\Delta_{t},q_{t})$) and $\sigma_{x}$, $\sigma_{z}$ are the Pauli-X-operator and Pauli-Z-operator, respectively. The energy eigenvalues are given by $\epsilon_{\pm}=\pm\omega_{t}/2$, where we define the energy difference as $\omega_{t}:=\sqrt{\Delta_{t}^{2}+q_{t}^{2}}$. The energy eigenstates are given by $|\epsilon_{+}\rangle=\cos\theta_{t}|e\rangle+\sin\theta_{t}|g\rangle$ and $|\epsilon_{-}\rangle=\sin\theta_{t}|e\rangle-\cos\theta_{t}|g\rangle$, where $\theta_{t}:=(1/2)\text{cot}^{-1}(q_{t}/\Delta_{t})$, and $|e\rangle$ and $|g\rangle$ are the eigenstates of $\sigma_{z}$, i.e., $\sigma_{z}|e\rangle=|e\rangle$ and $\sigma_{z}|g\rangle=-|g\rangle$. The time-dependent Hamiltonian (38) can be realized, for example, in the charge qubit configuration, where $q_{t}$ can be varied by changing the applied magnetic flux, while $\Delta_{t}$ can be varied by changing the applied gate voltage Pekola et al. (2019).
@@
$\dot{\rho}_{t}=-i[H_{t},\rho_{t}]+\gamma_{\downarrow}\left(\sigma_{t}^{-}\rho\sigma_{t}^{+}-\frac{1}{2}\left\{\sigma_{t}^{+}\sigma_{t}^{-},\rho_{t}\right\}\right)+\gamma_{\uparrow}\left(\sigma_{t}^{+}\rho\sigma_{t}^{-}-\frac{1}{2}\left\{\sigma_{t}^{-}\sigma_{t}^{+},\rho_{t}\right\}\right),$ (39)
where $\sigma_{t}^{+}=\left.\left|\epsilon_{+}\right\rangle\right.\left.\left\langle\epsilon_{-}\right.\right|$ and $\sigma_{t}^{-}=\left.\left|\epsilon_{-}\right\rangle\right.\left.\left\langle\epsilon_{+}\right.\right|$ are the raising and lowering operators between energy eigenstates, and $\gamma_{\downarrow}:=\gamma_{0}(N(\omega_{t})+1)$ and $\gamma_{\uparrow}:=\gamma_{0}N(\omega_{t})$ with $N(\omega_{t}):=1/(e^{\beta\omega_{t}}-1)$ are the transition rates satisfying the detailed balance condition $\gamma_{\downarrow}/\gamma_{\uparrow}=e^{\beta\omega_{t}}$. In the following, we use the notation $p_{\pm}=\langle\epsilon_{\pm}|\rho|\epsilon_{\pm}\rangle$ and $\rho_{+-}=\langle\epsilon_{+}|\rho|\epsilon_{-}\rangle$,
In this setup, we can derive a simplified analytical expression for the obtained bound (34). In particular, the counter-diabatic Hamiltonian is given by
and its variance reads
$V_{\rho_{t}}(H_{\rm cd})=\dot{\theta}^{2}_{t}\left(1-4(\Im[\rho_{+-}])^{2}\right).$ (41)
In addition, the activity-like quantity reads
$A(\rho_{t}^{\rm cl})=2\left(\frac{\dot{\omega}_{t}}{\gamma_{\downarrow}+\gamma_{\uparrow}}\right)^{2}A_{\rm act},$ (42)
where
$A_{\rm act}=\gamma_{\uparrow}p_{-}+\gamma_{\downarrow}p_{+}$ (43)
is the activity. This expression demonstrates that $A(\rho_{t}^{\rm cl})$ is essentially given by the total number of jump rates $A_{\rm act}$ multiplied by the square of the relative timescales between the change of the Hamiltonian $\dot{\omega}_{t}$ and the relaxation rates $\gamma_{\uparrow}+\gamma_{\downarrow}$.
@@ VII Conclusion
An interesting future direction is to further investigate the classical-quantum decomposition of non-adiabatic work and analyze the role of quantum coherence in thermodynamics. In the quasi-adiabatic regime, it has been reported in Ref. Miller et al. (2019) that generation of quantum coherence violates the fluctuation-dissipation relation. It has also been observed in various models that generation of quantum coherence is typically detrimental to the performance of heat engines Pekola et al. (2019). However, a general understanding on the effect of having coherence in the non-adiabatic work extraction processes is not complete and deserves further studies. Understanding the role of coherence in thermodynamics is anticipated to lead to obtaining the design principle of fast and efficient energy conversion and information processing devices.
It is also interesting to investigate further the geometric and thermodynamic properties of the obtained bound for nonequilibrium states. Extending the concepts of geometric quantities defined for adiabatic or equilibrium states and their optimization methods Brandner and Saito (2020); Scandi and Perarnau-Llobet (2019); Abiuso et al. (2020); Miller et al. (2019); Rolandi et al. (2023) to nonequilibrium setting is another important future direction. We also note that the obtained thermodynamic speed limit bounds the absolute value of the non-adiabatic work, while the sign of the non-adiabatic work is shown to be non-negative in the slow-driving regime Miller et al. (2019). It would be interesting if we can further generalize our approach and obtain bounds beyond its absolute value. It would also be an interesting direction if we could derive thermodynamic uncertainty relation bounds on non-adiabatic work, for example, utilizing the fluctuation theorem uncertainty relations Timpanaro et al. (2019); Hasegawa and Van Vu (2019b).
@@ Appendix A Properties of the Drazin pseudo inverse and its adjoint
Next, let us consider the Drazin inverse of $\mathcal{H}_{t}$ (2). We first note that the steady state of $\mathcal{H}_{t}$ is no longer unique, since any $\rho_{t}^{\rm cl}$ satisfies $\mathcal{H}_{t}[\rho_{t}^{\rm cl}]=0$. For given $\rho_{t}$, the steady-state of $\mathcal{H}_{t}$ reads
$\lim_{\nu\rightarrow\infty}e^{\nu\mathcal{H}_{t}}\rho_{t}=\sum_{n}\langle\epsilon_{n}|\rho_{t}|\epsilon_{n}\rangle|\epsilon_{n}\rangle\langle\epsilon_{n}|=\rho_{t}^{\rm cl},$ (46)
@@
$\mathcal{H}_{t}^{+}[\rho_{t}^{\rm cl}]=\mathrm{Tr}\left[\mathcal{H}_{t}^{+}[X]\right]=0.$ (49)